# ELMS: Elasticized Large Language Models On Mobile Devices

### v1

@@ 3.1. Overview
Design goal. ELMS aims to provide LLM service that adapts to a specific Service Level Objective (SLO) of resource constraint per request (prompt), while maximizing the service quality (i.e., LLM generation accuracy).
SLO definition. In this paper, we define SLO of LLM service as a tuple $<\zeta_{TTFT},\zeta_{TPOT}>$, where $\zeta$ is the compression ratio to full LLMaaS latency. The SLOs that an LLMaaS should serve is pre-defined by the service developers.
Workflow. ELMS’s idea is to orchestrate model- and prompt-level elastification, both of which need offline preparation. To achieve this, its workflow is designed as a cloud, offline stage and a device, online stage, as shown in Figure 6.
At cloud offline stage, on one hand, the model is elasticized to various levels of sub-models that share the memory and can be cost-effectively switched (§$\S$3.2). On the other hand, we on-cloud fine-tune a TLM for prompt elastification (§$\S$3.3).
At device online stage, the elasticized LLM and fine-tuned TLM are deployed on mobile devices as a service. For each LLM service request, the prompt and the corresponding SLO are fed into the fine-tuned TLM. The TLM then outputs a compressed prompt and selects a sub-model with proper size. Finally, an LLM inference that satisfies the given SLO is performed atop the sub-model and the compressed prompt.
Figure 6. Workflow of ELMS.
@@ 3.2. Model elastification
This property indicates that a dense operator kernel can equivalently process these units in arbitrary order without any runtime data re-layout. The rationale behind it is that the Reduce operator (e.g., sum/min/max) satisfies the commutative and associative laws. A basic block that contains such units is $y=xW_{1}W_{2}$ in Figure 7. Its permutation consistent unit is a column of $W_{1}$ together with the corresponding row of $W_{2}$. If we permute the weights as shown in Figure 7, the intermediate activation $xW^{\prime}_{1}$ will be permuted in response. Nevertheless, $W^{\prime}_{2}$ is also permuted in the same order, so the multiplication of MatMul operator can still be performed correctly. Since the following addition of MatMul operator is a Reduce operator, the calculated $x{W^{\prime}}_{1}{W^{\prime}}_{2}$ is exactly the same as $xW_{1}W_{2}$. Notably, different from prior work (Zheng et al., 2023b) that also leverages the Reduce operator to permute DNNs, ELMS’s key insight is to identify such a joint unit in two-layer blocks that are ubiquitous in Transformers. Permutation of this unit can be made completely offline, while (Zheng et al., 2023b) still needs online reordering the input with a single operator level abstraction.
@@
Our method: one-shot unit-reordering. Based on Property 1 and Property 2, we propose a novel model elastification method as shown in Figure 9. Its key idea is to “atomize” the Transformer into the units shown in Figure 8, and then group them to construct a series of sub-models that each is contiguous in memory.
$\bullet$ Offline. Specifically, ELMS first profiles importance of each unit offline (detailed later). Since these units are permutation consistent, ELMS freely reorders them by their importance in descending order (if importance is the higher the better), starting from the base address of the weight tensor. Notably, the reordering is only performed intra-block, e.g., reordering the unit of attention heads in the same attention block. Then, ELMS groups these units into sub-models. For example, in Figure 9, sub-model in $level^{M}_{3}$ contains units with indices (not address) “1 5 8 2”, and sub-model in $level^{M}_{1}$ contains “1 5 8 2 3 4 7”. The sub-model sizes and numbers $\{level^{M}_{1},\cdots,level^{M}_{J}\}$ are pre-defined by the LLMaaS developers. In practice, we set this to a fine enough granularity (a global ratio of 20% to 100% in a step size of 10%, by default). Such a ratio is evenly shared by all Transformer layers that need elastification. After that, low-rank adapters (Hu et al., 2021) are attached to each sub-model to recover potential generation quality loss (if there is any) of these sub-models. We elaborate such a recovery process in the following parts. So far, the LLM weights have been elasticized into a series of high-quality sub-models that run in various speed. We demonstrate the quality of generated sub-models in Figure 10.
$\bullet$ Online. In Figure 9b, the upgrading of model is performed in the following steps: ELMS first detaches the corresponding adapter from sub-model $level^{M}_{1}$, which has served the last request. Then, it moves the ending memory pointer of the weights tensor from the address of unit with index “2” to “7”. After that, ELMS attaches another adapter to sub-model $level^{M}_{3}$, and the upgrading is finished. Such a process is very cost-effective on mobile devices — it does not involve any data movement compared to traditional pruning methods, and can still utilize deeply-optimized dense kernels provided by NN libraries. For instance, upgrading $W_{Q}$ to 4096$\times$4096 size only takes 2 ms on Redmi K60 Champion smartphone, while a naive pruning method must undergo a 140ms data movement.
Figure 9. ELMS offline profiles and reorders permutation consistent units in one-shot; it online switches (i.e., upgrade/downgrade) sub-models by cost-effectively attaching/detaching the corresponding adapters and moving the memory pointer.
@@
Figure 10. (a) ELMS generates sub-models that with consistently higher quality than pruning and pruning+recovery (Ma et al., 2023). (b) A small portion of layers. i.e., “anchor layers”, is much more important than others.
Profiling unit importance through explainable-AI. Parameters of neuron networks are known to feature diverse importance. For instance, a weight element with higher magnitude may contribute more to NN capability (Han et al., 2015; Janowsky, 1989; Li et al., 2017). Inspired by the concept of eXplainable-AI (XAI) (Huang et al., 2023; Selvaraju et al., 2019; Sundararajan et al., 2017), ELMS profiles unit importance with a more accurate method, i.e., by the next-token prediction loss function $L$ on a calibration corpus $C$. The intuition behind XAI is that, if a unit is more important, it should make $L$ larger on $C$ when been removed. Specifically, we define importance of unit $i$ as $imp_{i}=|L-L_{W_{i}=0}|$. By Maclaurin series (Swokowski, 1979), we get
(2) $\begin{split}imp_{i}=|L-L_{W_{i}=0}|=|\frac{\partial L}{\partial W_{i}}W_{i}+o(||W_{i}||^{2})|.\end{split}$
Since the second term is a higher-order infinitesimal, ELMS then takes $|\frac{\partial L}{\partial W_{i}}W_{i}|$ as a estimation of unit importance. By default, $C$ is set to a sub-set of bookcorpus (Zhu et al., 2015), which is a general-purpose language-modeling dataset.
@@
Task-agnostic low-rank recovery of sub-models. We add Low-Rank Adapters (LoRAs) (Hu et al., 2021) to the frozen $W_{Q/K/V/O}$ and $W_{up/down}$ of each sub-model to recover them from potential accuracy loss. A LoRA is two low-rank matrices $A\in\mathcal{R}^{n\times r}$ and $B\in\mathcal{R}^{r\times m}$ that trained as a side branch of main weights. ELMS’s default setting of $r$ is 8, an empirically common practice in LoRA-based tunings. LoRA weights are only 0.1%–0.5% of the entire LLM weights, and thus such a method only introduces ¡5% extra memory overhead even under ELMS’s fine-grained sub-model settings. The switching overhead is also minimized since the attaching/detaching operations are all low-rank MatMul and element-wise MatAdd.
Different from traditional language models that need tuning on specific downstream tasks (e.g., Bert (Devlin et al., 2019), T5 (Raffel et al., 2023)), LLM commonly serves as a generic task solver. Thereby, ELMS’s sub-model recovery is task-agnostic. LoRAs are trained with next-token prediction loss that identical to the pre-training process. Thanks to LoRA’s preservation of the LLM backbone’s capability, a general, high-quality and moderate-size corpus can handle this recovery well. By default, each sub-model is recovered on about 50M tokens of Alpaca-cleaned (Taori et al., 2023; alp, 2024) dataset. We also discuss the impact of recovery data in §$\S$5.5.
Remarks ELMS’s model elastification generates fine-grained, high-quality sub-models with both acceptable offline overhead and negligible online overhead.
@@ 3.3. Prompt elastification
Dual-head TLM. ELMS tackles the challenges mentioned in §$\S$2.2 by a dual-head Tiny Language Model (TLM). As shown in Figure 11, the TLM is a variant of mobile-friendly pre-trained language model MobileBert (Sun et al., 2020), a compact model with only 20% parameters of BERT_base (Devlin et al., 2019) yet just 0.7% accuracy loss on GLUE benchmark (Wang et al., 2019). We make the following modifications. Firstly, the SLO of the current request is marked in natural language and inserted into the embedding layer of MobileBert as special tokens. For instance, “[05]” represents the prefill SLO is 50% TTFT; “¡08¿” is 80% TPOT. These special tokens are initialized to word vectors that orthogonal to each other. Secondly, the TLM is designed with two separate heads, named score-head and decision-head. The score-head treats each token of the prompt as a two-class classification problem, where each token can be classified into “discard” or “retain”. The decision-head treats the entire sequence of prompt with SLO as two multi-class classification problems. Each possible model elastification level (i.e., the sub-model size, discussed in §$\S$3.2) is a class of one problem; possible prompt elastification levels are classes of the other problem. Akin to model elastification, the prompt is also pre-defined to multiple fine-grained levels that denoted $\{level^{P}_{1},\cdots,level^{P}_{K}\}$ by the LLMaaS developers. By default, we set it in alignment with model elastification levels. Besides, the two heads share the same bottom layers (12 out of 24 layers by default) based on the rationale that bottom layers of DNNs are mainly responsible for capturing basic instead of task-specific information. In doing so, the overhead of TLM inference/training is further minimized.
TLM inference. At decision-head, ELMS takes the class with the max probability as its decision. For example, in Figure 11, the decision-head selects a 60% prompt and 80% model. At score-head, ELMS ranks the prompt tokens by the probability of “retain”, then selects top ones according to decision-head. Notably, if the TLM outputs a decision that cannot meet the SLO (which is nearly yet not impossible due to the black-box property of DNNs), ELMS will execute a runtime check and the decision will fall back to a random one that stringently meet the SLO. After inference, we get a compressed prompt and a selected sub-model.
@@
$\bullet$ Decision-head. After training score-head, we train decision-head with the assistance of both the elasticized LLM and the score-head. The intuition is that we can traverse the decision space and learn the optimal one offline. Specifically, the training data is collected by a self-induced labelling process. In Figure 12, we provide an illustrative example. For each prompt and SLO (pre-defined by LLMaaS developer, see §$\S$3.1), we enumerate all its possible [^6] The LLMaaS developers pre-execute a one-shot profiling on a testing device within 1 hour under the guidance of Formula 1 and leverage the profiled data to judge whether a decision can meet a SLO. decisions, and label it with the most lightweight one under which the LLM can output correct answer. If all possible decisions fail, the decision is assigned to a default one (e.g., random). By default, ELMS collects training samples from a comprehensive benchmark MMLU-Pro (Wang et al., 2024), which contains 17 domains of question-answer pairs. In Figure 13, the decision head demonstrates significantly higher quality than random decision, and approaches oracle. We also discuss the impact of TLM training data in §$\S$5.5.
Remarks ELMS’s prompt elastification efficiently and effectively elasticizes prompt and orchestrates model- and prompt-dimension of elastic LLMaaS.
@@ 4. Implementation
Offline stage. We build the elasticized LLM and the TLM on top of Pytorch (Paszke et al., 2019). We modify the modeling.py of Huggingface Transformers [^7] Commit hash c31473ed4492fdf26aec4173451f31590021862f (Wolf et al., 2020) library to identify the permutation consistent units and profile their importance. In doing so, ELMS can easily be made forward-compatible with LLMs released in the future. The offline elastification is performed on a cloud server equipped with 8 NVIDIA A40 GPUs.
On-device LLMaaS. We build an LLM service on top of mllm [^8] Commit hash ed766fae54d9f1bf2b6b25018e6cc434ce223303 (mll, 2024), which is a lightweight yet powerful on-device LLM inference library written in pure C++ and assembly language. We pack the LLM inference program as a standalone binary file and run it as an independent process. The apps interact with it through interfaces like bindLLMService() and callLLM(). To facilitate the upgrade/downgrade of elasticized sub-models without incurring inference-time performance degradation, the Linear is replaced by ElasticLinear. Specifically, we wrap the original dense kernel with an additional memory pointer that specifies the addresses of sub-model weights. We optimize the low-rank MatMul and MatAdd of LoRA with ARM NEON (arm, 2024). The LLMaaS runs on Commercial Off-The-Shelf (COTS) smartphones.
@@ 5.1. Experimental settings
Testbed. We conduct experiments on the following testbeds. On cloud, we use a server with a 64-core CPU (750GB RAM) and 8 A40 GPUs (45GB HBM each). On device, we test ELMS across COTS smartphones listed in Table 2.
@@
SLOs. We randomly set 6 SLOs based on a stepwise sensitivity hierarchy as shown in Table 3. We also enumerate more SLOs in Figure 19 as supplementary experiments.
LLMaaS workload. We evaluate ELMS on both standalone datasets and end-to-end synthesized traces.
@@
GMail ¡80%, 90%¿ ARC_E Gboard ¡20%, 60%¿ OBQA
Octopus ¡60%, 80%¿ Octopus XiaoAi ¡20%, 50%¿ ARC_E
@@
ELMS configurations. The training/fine-tuning/calibration data, elastification levels and TLM configuration are the same as those described in §$\S$3.
Baselines. We compare ELMS to the following alternatives. (1) Directly employing pre-trained from scratch LLMs (PFS) for diverse SLOs is a strong yet plausible baseline. Due to the unaffordable GPU resource consumption, we select the off-the-shelf OPT (Zhang et al., 2022) family. As far as we know, it provides the richest variants with fewer than 7B parameters (5 models from 125M to 6.7B). (2) LLMPruner (Ma et al., 2023) (LPruner) is a State-of-The-Art (SoTA) parameter pruning method for elastification. (3) Layer-wise elastification (lay, 2024) (LE) prunes parameters at layer level. (4) LLMLingua2 (Pan et al., 2024) + Contextual sparsity (Liu et al., 2023) (LG2+CS) compresses prompts at prefill stage, and dynamically activates MLP neurons for each token at decode stage.

### v2

@@ 3.1. Overview
Design goal. ElastiLM aims to provide LLM service that adapts to a specific Service Level Objective (SLO) of resource constraint per request (prompt), while maximizing the service quality (i.e., LLM generation accuracy).
SLO definition. In this paper, we define SLO of LLM service as a tuple $\zeta_{TTFT},\zeta_{TPOT}>$, where $\zeta$ is the compression ratio to full LLM latency. The SLOs that an LLM service should serve is pre-defined by the service developers.
Workflow. As shown in Figure 6, ElastiLM features a two-stage workflow. At cloud offline stage, on one hand, the model is elasticized to various levels of sub-models that share the memory and can be cost-effectively switched ($\S$3.2). On the other hand, we on-cloud fine-tune a TLM for prompt elastification ($\S$3.3). At device online stage, the elasticized LLM and fine-tuned TLM are deployed on mobile devices as a service. For each LLM service request, the prompt and the corresponding SLO are fed into the fine-tuned TLM. The TLM then outputs a compressed prompt and selects a sub-model with proper size. Finally, an LLM inference that satisfies the given SLO is performed atop the sub-model and the compressed prompt.
Figure 6. Workflow of ElastiLM.
@@ 3.2. Model elastification
This property indicates that a dense operator kernel can equivalently process these units in arbitrary order without any runtime data re-layout. The rationale behind it is that the Reduce operator (e.g., sum/min/max) satisfies the commutative and associative laws. A basic block that contains such units is $y=xW_{1}W_{2}$ in Figure 7. Its permutation consistent unit is a column of $W_{1}$ together with the corresponding row of $W_{2}$. If we permute the weights as shown in Figure 7, the intermediate activation $xW^{\prime}_{1}$ will be permuted in response. Nevertheless, $W^{\prime}_{2}$ is also permuted in the same order, so the multiplication of MatMul operator can still be performed correctly. Since the following addition of MatMul operator is a Reduce operator, the calculated $x{W^{\prime}}_{1}{W^{\prime}}_{2}$ is exactly the same as $xW_{1}W_{2}$. Notably, different from prior work (Zheng et al., 2023b) that also leverages the Reduce operator to permute DNNs, ElastiLM’s key insight is to identify such a joint unit in two-layer blocks that are ubiquitous in Transformers. Permutation of this unit can be made completely offline, while (Zheng et al., 2023b) still needs online reordering the input with a single operator level abstraction.
@@
Our method: one-shot unit-reordering. Based on Property 1 and Property 2, we propose a novel model elastification method as shown in Figure 9. Its key idea is to “atomize” the Transformer into the units shown in Figure 8, and then group them to construct a series of sub-models that each is contiguous in memory. To illustrate, basically, in Figure 7 each column in $W_{1}$ and the row with same index in $W_{2}$ will be assigned with an importance score offline. Then the columns/rows are reordered in $W_{1}$$W_{2}$ to make a submodel contiguous. At online, $W^{\prime}_{1}$ and $W^{\prime}_{2}$ only need to slice a submatrix (e.g., with indices 1,4 out of $W^{\prime}_{1}$$W^{\prime}_{2}$) out of the original one by a zero-cost movement of memory pointer. Detailed later.
$\bullet$ Offline. Specifically, ElastiLM first profiles importance of each unit offline (detailed later). Since these units are permutation consistent, ElastiLM freely reorders them by their importance in descending order (if importance is the higher the better), starting from the base address of the weight tensor. Notably, the reordering is only performed intra-block, e.g., reordering the unit of attention heads in the same attention block. Then, ElastiLM groups these units into sub-models. For example, in Figure 9, sub-model in $level^{M}_{3}$ contains units with indices (not address) “1 5 8 2”, and sub-model in $level^{M}_{1}$ contains “1 5 8 2 3 4 7”. The sub-model sizes and numbers $\{level^{M}_{1},\cdots,level^{M}_{J}\}$ are pre-defined by the developers. In practice, we set this to a fine enough granularity (a global ratio of 20% to 100% in a step size of 10%, by default). Such a ratio is evenly shared by all Transformer layers that need elastification. After that, low-rank adapters (Hu et al., 2021) are attached to each sub-model to recover potential generation quality loss (if there is any) of these sub-models. We elaborate such a recovery process in the following parts. So far, the LLM weights have been elasticized into a series of high-quality sub-models that run in various speed. We demonstrate the quality of generated sub-models in Figure 10.
$\bullet$ Online. In Figure 9b, the upgrading of model is performed in the following steps: ElastiLM first detaches the corresponding adapter from sub-model $level^{M}_{1}$, which has served the last request. Then, it moves the ending memory pointer of the weights tensor from the address of unit with index “2” to “7”. After that, ElastiLM attaches another adapter to sub-model $level^{M}_{3}$, and the upgrading is finished. Such a process is very cost-effective on mobile devices — it does not involve any data movement compared to traditional pruning methods, and can still utilize deeply-optimized dense kernels provided by NN libraries. For instance, upgrading $W_{Q}$ to $\times4096$ size only takes 2 ms on Redmi K60 Champion smartphone, while a naive pruning method must undergo a 140ms data movement.
Figure 9. ElastiLM offline profiles and reorders permutation consistent units in one-shot; it online switches (i.e., upgrade/downgrade) sub-models by cost-effectively attaching/detaching the corresponding adapters and moving the memory pointer.
@@
Figure 10. (a) ElastiLM generates sub-models that with consistently higher quality than pruning and pruning+recovery (Ma et al., 2023). (b) A small portion of layers. i.e., “anchor layers”, is much more important than others.
Profiling unit importance through explainable-AI. Parameters of neuron networks are known to feature diverse importance. For instance, a weight element with higher magnitude may contribute more to NN capability (Han et al., 2015; Janowsky, 1989; Li et al., 2017). Inspired by the concept of eXplainable-AI (XAI) (Huang et al., 2023; Selvaraju et al., 2019; Sundararajan et al., 2017), ElastiLM profiles unit importance with a more accurate method, i.e., by the next-token prediction loss function $L$ on a calibration corpus $C$. The intuition behind XAI is that, if a unit is more important, it should make $L$ larger on $C$ when been removed. Specifically, we define importance of unit $i$ as $imp_{i}=|L-L_{W_{i}=0}|$. By Maclaurin series (Swokowski, 1979), we get
[…]
Since the second term is a higher-order infinitesimal, ElastiLM then takes $|\frac{\partial L}{\partial W_{i}}W_{i}|$ as a estimation of unit importance. By default, $C$ is set to a sub-set of bookcorpus (Zhu et al., 2015), which is a general-purpose language-modeling dataset.
@@
Task-agnostic low-rank recovery of sub-models. We add Low-Rank Adapters (LoRAs) (Hu et al., 2021) to the frozen $W_{Q/K/V/O}$ and $W_{up/down}$ of each sub-model to recover them from potential accuracy loss. A LoRA is two low-rank matrices $A\in\mathcal{R}^{n\times r}$ and $B\in\mathcal{R}^{r\times m}$ that trained as a side branch of main weights. ElastiLM’s default setting of $r$ is 8, an empirically common practice in LoRA-based tunings. LoRA weights are only 0.1%–0.5% of the entire LLM weights, and thus such a method only introduces ¡5% extra memory overhead even under ElastiLM’s fine-grained sub-model settings. The switching overhead is also minimized since the attaching/detaching operations are all low-rank MatMul and element-wise MatAdd.
Different from traditional language models that need tuning on specific downstream tasks (e.g., Bert (Devlin et al., 2019), T5 (Raffel et al., 2023)), LLM commonly serves as a generic task solver. Thereby, ElastiLM’s sub-model recovery is task-agnostic. LoRAs are trained with next-token prediction loss that identical to the pre-training process. Thanks to LoRA’s preservation of the LLM backbone’s capability, a general, high-quality and moderate-size corpus can handle this recovery well. By default, each sub-model is recovered on about 50M tokens of Alpaca-cleaned (Taori et al., 2023; alp, 2024) dataset. We also discuss the impact of recovery data in $\S$$5.$
Remarks ElastiLM’s model elastification generates fine-grained, high-quality sub-models with both acceptable offline overhead and negligible online overhead.
@@ 3.3. Prompt elastification
Dual-head TLM. ElastiLM tackles the challenges mentioned in $\S$2.2 by a dual-head Tiny Language Model (TLM). As shown in Figure 11, the TLM is a variant of mobile-friendly pre-trained language model MobileBert (Sun et al., 2020), a compact model with only 20% parameters of BERT_base (Devlin et al., 2019) yet just 0.7% accuracy loss on GLUE benchmark (Wang et al., 2019). We make the following modifications. Firstly, the SLO of the current request is marked in natural language and inserted into the embedding layer of MobileBert as special tokens. For instance, “[05]” represents the prefill SLO is 50% TTFT; “¡08¿” is 80% TPOT. These special tokens are initialized to word vectors that orthogonal to each other. Secondly, the TLM is designed with two separate heads, named score-head and decision-head. The score-head treats each token of the prompt as a two-class classification problem, where each token can be classified into “discard” or “retain”. The decision-head treats the entire sequence of prompt with SLO as two multi-class classification problems. Each possible model elastification level (i.e., the sub-model size, discussed in $\S$3.2) is a class of one problem; possible prompt elastification levels are classes of the other problem. Akin to model elastification, the prompt is also pre-defined to multiple fine-grained levels $\{level^{P}_{1},\cdots,level^{P}_{K}\}$ by the developers. By default, we set it in alignment with model elastification levels. Besides, the two heads share the same bottom layers (12 out of 24 layers by default) based on the rationale that bottom layers of DNNs are mainly responsible for capturing basic instead of task-specific information. In doing so, the overhead of TLM inference/training is further minimized.
TLM inference. At decision-head, ElastiLM takes the class with the max probability as its decision. For example, in Figure 11, the decision-head selects a 60% prompt and 80% model. At score-head, ElastiLM ranks the prompt tokens by the probability of “retain”, then selects top ones according to decision-head. For example, in Figure 11, the scores are ranked by the green probabilities. Notably, if the TLM outputs a decision that cannot meet the SLO (which is nearly yet not impossible due to the black-box property of DNNs), ElastiLM will execute a runtime check and the decision will fall back to a random one that stringently meet the SLO. After inference, we get a compressed prompt and a selected sub-model.
@@
$\bullet$ Decision-head. After training score-head, we train decision-head with the assistance of both the elasticized LLM and the score-head. The intuition is that we can traverse the decision space and learn the optimal one offline. Specifically, the training data is collected by a self-induced labelling process. In Figure 12, we provide an illustrative example. For each prompt and SLO (pre-defined by developer, see $\S$3.1), we enumerate all its possible [^6] The developers pre-execute a one-shot profiling on a testing device within 1 hour under the guidance of Formula 1 and leverage the profiled data to judge whether a decision can meet a SLO. decisions, and label it with the most lightweight one under which the LLM can output correct answer. If all possible decisions fail, the decision is assigned to a default one (e.g., random). By default, ElastiLM collects training samples from a comprehensive benchmark MMLU-Pro (Wang et al., 2024), which contains 17 domains of question-answer pairs. In Figure 13, the decision head demonstrates significantly higher quality than random decision, and approaches oracle. We also discuss the impact of TLM training data in $\S$$5.$
Remarks ElastiLM’s prompt elastification efficiently and effectively elasticizes prompt and orchestrates model- and prompt-dimension of elastic LLM service.
@@ 4. Implementation
Offline stage. We build the elasticized LLM and the TLM on top of Pytorch (Paszke et al., 2019). We modify the modeling.py of Huggingface Transformers [^7] Commit hash c31473ed4492fdf26aec4173451f31590021862f (Wolf et al., 2020) library to identify the permutation consistent units and profile their importance. In doing so, ElastiLM can easily be made forward-compatible with LLMs released in the future. The offline elastification is performed on a cloud server equipped with 8 NVIDIA A40 GPUs.
On-device stage. We build an LLM service on top of mllm [^8] Commit hash ed766fae54d9f1bf2b6b25018e6cc434ce223303 (mll, 2024), which is a lightweight yet powerful on-device LLM inference library written in pure C++ and assembly language. We pack the LLM inference program as a standalone binary file and run it as an independent process. The apps interact with it through interfaces like bindLLMService() and callLLM(). To facilitate the upgrade/downgrade of elasticized sub-models without incurring inference-time performance degradation, the Linear is replaced by ElasticLinear. Specifically, we wrap the original dense kernel with an additional memory pointer that specifies the addresses of sub-model weights. We optimize the low-rank MatMul and MatAdd of LoRA with ARM NEON (arm, 2024). The LLM service runs on Commercial Off-The-Shelf (COTS) smartphones.
@@ 5.1. Experimental settings
Testbed. We conduct experiments on the following testbeds. On cloud, we use a server with a 64-core CPU (750GB RAM) and 8 A40 GPUs (45GB HBM each). On device, we test ElastiLM across COTS smartphones listed in Table 2.
@@
SLOs. We randomly set 6 SLOs based on a stepwise sensitivity hierarchy as shown in Table 3. We also enumerate more varying SLOs in Figure 18 for a real-world discussion. Rewind is a content comprehension app, which requires relaxed SLO; GMail is a natural language summary and reasoning task, which requires relatively relaxed SLO; Octopus and Shortcuts are on-device agent tasks, which need moderate SLO as they involve interactions with the environment; Gboard and XiaoAi are AI assistants, which need real-time response, i.e., a relatively tight SLO.
Workload. We evaluate ElastiLM on both standalone datasets and end-to-end synthesized traces.
@@
GMail ¡80%, 90%¿ ARC_E Gboard ¡20%, 60%¿ PIQA
Octopus ¡60%, 80%¿ Octopus XiaoAi ¡20%, 50%¿ SCIQ
@@
ElastiLM configurations. The training/fine-tuning/calibration data, elastification levels and TLM configuration are the same as those described in $\S$3.
Baselines. We compare ElastiLM to the following alternatives. (1) Directly employing pre-trained from scratch LLMs (PFS) for diverse SLOs is a strong yet plausible baseline. Due to the unaffordable GPU resource consumption, we select the off-the-shelf OPT (Zhang et al., 2022) family. As far as we know, it provides the richest variants with fewer than 7B parameters (5 models from 125M to 6.7B). (2) LLMPruner (Ma et al., 2023) (LPruner) is a State-of-The-Art (SoTA) parameter pruning method for elastification. (3) Layer-wise elastification (lay, 2024) (LE) prunes parameters at layer level. (4) LLMLingua2 (Pan et al., 2024) + Contextual sparsity (Liu et al., 2023) (LG2+CS) compresses prompts at prefill stage, and dynamically activates MLP neurons for each token at decode stage. (5) Other strong pruning baselines: LaCo (Yang et al., 2024a), ShortGPT (Men et al., 2024) and AttnDrop (He et al., 2024).